What is the steric number for the three indicated carbon atoms (C1,C2,and C3) in the following molecule? 1st attempt Feedback See Periodic Ta VIEW SOLUT 01/10 /> O OF 10 QUESTIONS COMPLETED OType here to search
media%2F872%2F8727d9d4-ab68-4360-afaf-24
Answer #1

Steric Number = The no atoms bonded to the central atom + lone pair of electrons

Or simply no of sigma bonds + no of lone pairs

In our example for C1 = 4 ( 3 sigma bonds with 3H and one sigma bond with C)

for C2 = 3 (2 sigma bonds with 2 Carbons + 1 sigma bond with Oxygen )

for C3 = 2 ( 1sigma bond with carbonyl carbon + 1 sigma bond with caynide carbon)